A photon and an electron each have an energy of
6×10^3.
What are their
wavelengths? Which of these would you use to probe atomic structures?
1
Expert's answer
2020-03-17T10:09:17-0400

"E=6\cdot10^3" "eV"


"\lambda=\frac{h}{\sqrt{2mE}}"


For electron


"\lambda_e=\frac{h}{\sqrt{2mE}}=\frac{6.62\cdot10^{-34}}{\sqrt{2\cdot 9.1\cdot10^{-31}\cdot 1.6\cdot10^{-19}\cdot 6000}}=1.58\cdot10^{-11} m=15.8" "pm"


For proton


"\lambda_p=\frac{h}{\sqrt{2mE}}=\frac{6.62\cdot10^{-34}}{\sqrt{2\cdot 1.67\cdot10^{-27}\cdot 1.6\cdot10^{-19}\cdot 6000}}=1.17\cdot10^{-12} m=1.17" "pm"



I would use electrons to probe atomic structures.